Determine molecular mass and uncertainty for C6H5OH. Take sig figs into account.
C = 12.0107 ± 8
H = 1.00794 ± 7
O = 15.9994 ± 3
1
Expert's answer
2013-02-05T11:19:55-0500
Mw of C6H5OH = 6*C + 6*H + O = 6*(12.0107± 8) +6*(1.00794± 7)+15.9994 ± 3 = (72,0642± 48) + (6,04764 ± 42) +(15.9994 ± 3)
= 94,11124 ± 552